OpenStudy (anonymous):

Jack has a fair coin. He tosses the coin four times. What is the probability that Jack gets heads, then heads, then tails, and then tails?

OpenStudy (perl):

The question is asking for P( H & H & T & T ) Each toss is independent , therefore P( H & H & T & T ) = P(H)* P(H) * P(T) * P (T)

OpenStudy (anonymous):

so...what do i do? multiply?

OpenStudy (anonymous):

@perl

jimthompson5910 (jim_thompson5910):

yes you multiply and P(Heads) = 1/2 P(Tails) = 1/2 because either side has the same chances of being landed on

OpenStudy (anonymous):

im asking, what do i multiply? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

P(H)* P(H) * P(T) * P (T) turns into (1/2)*(1/2)*(1/2)*(1/2) and then you evaluate that

OpenStudy (anonymous):

1/16

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (thomas5267):

BTW, a unique sequence of n fair coin toss has a probability of \(2^{-n}\).